# A Comparison Between Invariant and Equivariant Classical and Quantum Graph Neural Networks

### v2

@@ 5 Conclusions
Through several computational experiments, the quantum GNNs seem to exhibit enhanced classifier performance over their classical GNN counterparts based on the best test AUC scores produced after the training of the models while relying on a similar number of parameters, hyperparameters, and model structures. These results seem promising for the quantum advantage over classical models. Despite this result, the quantum algorithms take over $100$ times as long to train than the classical networks. This may be attributed to the limited capabilities in the quantum APIs, where the inability to train with broadcastable unitary operators forced the quantum models to take in a batch size of one, or train on a single graph at a time.
The action of the equivariance in the EGNN and EQGNN could be further explored and developed. This is especially true in the quantum case where more general permutation and SU(2) equivariance have been explored Skolik et al. (2023); Mernyei et al. (2022); East et al. (2023); Zheng et al. (2023). Expanding the flexibility of the networks to an arbitrary number of nodes per graph should also offer increased robustness, however, this may continue to pose challenges in the quantum case due to the current limited flexibility of quantum software. A variety of different forms of the networks can also be explored, such as introducing attention components and altering the structure of the quantum graph layers to achieve enhanced performance of both classical and quantum GNNs. In particular, one can further parameterize the quantum graph layer structure to better align with the total number of parameters used in the classical structures. These avenues will be explored in future work.
@@ 6 Software and Code
Acknowledgements
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0025759. SG is supported in part by the U.S. Department of Energy (DOE) under Award No. DE-SC0012447. KM is supported in part by the U.S. DOE award number DE-SC0022148. KK is supported in part by US DOE DE-SC0024407. CD is supported in part by College of Liberal Arts and Sciences Research Fund at the University of Kansas. CD, RF, EU, MCC and TM were participants in the 2023 Google Summer of Code.
@@ Appendix A Equivariant coordinate update function
Let $T_{g}:X\to X$ be the set of translational and rotational group transformations with elements $g\in T_{g}\subset G$ which act on the vector space $X$. The function $\varphi:X\to X$ defined by
@@
Let a general transformation $g\in T_{g}$ act on $X$ by $gX=RX+T$, where $R\in T_{g}$ denotes a general rotation and $T\in T_{g}$ a general translation. Then, under the transformation $g$ on $X$ of the function $\varphi$ defined above, we have
@@ Appendix B Coupling Hamiltonian simplification
The reduction of the coupling Hamiltonian becomes
@@
and multiplying on the left by $\hat{I}_{j}$ and on the right by $\hat{I}_{k}$ admits
@@ Appendix C Quantum product state permutation equivariance
For $V$ a commutable vector space, the product state $\bigotimes_{i=1}^{m}\mathbf{v}_{i}:V^{n}\times\cdots\times V^{n}\to V^{n^{m}}$ is permutation equivariant with respect to the sum of its entries. We prove the $n=2$ case for all $m\in\mathbb{Z}_{>0}$.
@@
(By Induction) Assume we have $n=1$ final qubit state,
@@
the sum of the product state elements will be trivially equivariant with respect to similar graphs. If we have $n=2$ final qubit states, the product state is
@@
for commutative spaces where $v_{i}^{j}\in\mathbb{C}$ and $\{1,2\},\{2,1\}$ should be regarded as ordered sets, which again shows the sum of the state elements remaining unchanged when the qubit states switch positions in the product. We will now assume the statement is true for $n=N$ final qubit states and proceed to show the $N+1$ case is true. The quantum product state over $N$ elements becomes
@@
which we assume to be permutation equivariant over the sum of its elements. We can rewrite the form of this state as
@@
where $A_{j}$ define the $2^{N}$ terms in the final product state. Replacing the $i+1$-th entry of the Kronecker product above with a new $N+1$-th state, we have
@@
When this occurs, this new state consisting of $2^{N+1}$ elements with the $N+1$ state in the $i+1$-th entry of the product can be written in terms of the old state with groupings of the new elements in $2^{N+1-i}$ batches of $2^{i}$ elements, i.e.
@@
which when summed becomes
@@
However, the $i+1$-th entry is arbitrary, and due to the summation permutation equivariance of the initial state $\bigotimes_{i=1}^{N}|\psi_{i}\rangle$, the sum $\sum_{j=1}^{2^{N}}A_{j}$ is equivariant, in fact invariant, under all reorderings of the elements $|\psi_{i}\rangle$ in the product $\bigotimes_{i=1}^{N}|\psi_{i}\rangle$. Therefore, we conclude $\bigotimes_{i=1}^{N+1}|\psi_{i}\rangle$ is permutation equivariant with respect to the sum of its elements. ∎
To show a simple illustration of why 39 is true, let us take two initial states and see what happens when we insert a new state between them, i.e. in the $2$nd entry in the product. This should lead to $42^{2+1-1}=2^{2}=$ 4 groupings of $22^{1}=$ 2 elements. To begin, we have
@@
and when we insert the new third state in the $1$-th entry of the product above, we have

### v3

@@ 5 Conclusions
Through several computational experiments, the quantum GNNs seemed to exhibit enhanced classifier performance compared with their classical GNN counterparts based on the best test AUC scores produced after the training of the models while relying on a similar number of parameters, hyperparameters, and model structures. These results seem promising for the quantum advantage over classical models. Despite this result, the quantum algorithms took over $100$ times longer to train than the classical networks. This was primarily due to the fact that we ran our quantum simulations on classical computers and not on actual quantum hardware. In addition, we were hindered by the limited capabilities in the quantum APIs, where the inability to train with broadcastable unitary operators forced the quantum models to take in a batch size of one or train on a single graph at a time.
The action of the equivariance in the EGNN and EQGNN could be further explored and developed. This is especially true in the quantum case where more general permutation and SU(2) equivariance have been explored Skolik et al. (2023); Mernyei et al. (2022); East et al. (2023); Zheng et al. (2023). Expanding the flexibility of the networks to an arbitrary number of nodes per graph should also offer increased robustness; however, this may continue to pose challenges in the quantum case due to the current limited flexibility of quantum software. A variety of different forms of the networks can also be explored. Potential ideas for this include introducing attention components and altering the structure of the quantum graph layers to achieve enhanced performance of both classical and quantum GNNs. In particular, one can further parameterize the quantum graph layer structure to better align with the total number of parameters used in the classical structures. These avenues will be explored in future work.
@@ 6 Software and Code
PyTorch and Pennylane were the primary APIs used in the formation and testing of these algorithms. The code corresponding to this study can be found at https://github.com/royforestano/2023_gsoc_ml4sci_qmlhep_gnn (accessed on 5 February 2024).
@@
Conceptualization, R.T.F.; methodology, M.C.C., G.R.D., Z.D., R.T.F., S.G., D.J., K.K., T.M., K.T.M., K.M. and E.B.U.; software, R.T.F.; validation, M.C.C., G.R.D., Z.D., R.T.F., T.M. and E.B.U.; formal analysis, R.T.F.; investigation, M.C.C., G.R.D., Z.D., R.T.F., T.M. and E.B.U.; resources, R.T.F., K.T.M. and K.M.; data curation, G.R.D., S.G. and T.M.; writing—original draft preparation, R.T.F.; writing—review and editing, S.G., D.J., K.K., K.T.M. and K.M.; visualization, R.T.F.; supervision, S.G., D.J., K.K., K.T.M. and K.M.; project administration, S.G., D.J., K.K., K.T.M. and K.M.; funding acquisition, S.G. All authors have read and agreed to the published version of the manuscript.
[…]
This study used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0025759. S.G. was supported in part by the U.S. Department of Energy (DOE) under Award No. DE-SC0012447. K.M. was supported in part by the U.S. DOE award number DE-SC0022148. K.K. was supported in part by US DOE DE-SC0024407. Z.D. was supported in part by College of Liberal Arts and Sciences Research Fund at the University of Kansas. Z.D., R.T.F., E.B.U., M.C.C., and T.M. were participants in the 2023 Google Summer of Code.
[…]
Not applicable.
[…]
The high-energy physics (HEP) dataset Pythia8 Quark and Gluon Jets for Energy Flow Komiske et al. (2019) was used in this analysis.
[…]
The authors declare no conflicts of interest. The funders had no role in the design of the study; in the collection, analyses, or interpretation of data; in the writing of the manuscript; or in the decision to publish the results.
[…]
Abbreviations The following abbreviations are used in this manuscript:
API Application Programming Interface
AUC Area Under the Curve
CNN Convolutional Neural Network
EGNN Equivariant Graph Neural Network
EQGNN Equivariant Quantum Graph Neural Network
FPR False Positive Rate
GAN Generative Adversarial Network
GNN Graph Neural Network
LHC Large Hadron Collider
MDPI Multidisciplinary Digital Publishing Institute
MLP Multilayer Perceptron
NLP Natural Language Processor
NN Neural Network
QGCNN Quantum Graph Convolutional Neural Network
QGNN Quantum Graph Neural Network
QGRNN Quantum Graph Recurrent Neural Network
TPR True Positive Rate
[…]
@@ Appendix A Equivariant Coordinate Update Function
Let $T_{g}:X\to X$ be the set of translational and rotational group transformations with elements $g\in T_{g}\subset G$ that act on the vector space $X$. The function $\varphi:X\to X$ defined by
@@
Let a general transformation $g\in T_{g}$ act on $X$ by $gX=RX+T$, where $R\in T_{g}$ denotes a general rotation, and $T\in T_{g}$ denotes a general translation. Then, under transformation $g$ on $X$ of function $\varphi$ defined above, we have
@@ Appendix B Coupling Hamiltonian Simplification
The reduction in the coupling Hamiltonian becomes
@@
and multiplying on the left by $\hat{I}_{j}$ and on the right by $\hat{I}_{k}$ produces
@@ Appendix C Quantum Product State Permutation Equivariance
For $V$, a commutable vector space, the product state $\bigotimes_{i=1}^{m}\mathbf{v}_{i}:V^{n}\times\cdots\times V^{n}\to V^{n^{m}}$ is permutation-equivariant with respect to the sum of its entries. We prove the $n=2$ case for all $m\in\mathbb{Z}_{>0}$.
@@
(By Induction) Assuming we have an $n=1$ final qubit state,
@@
the sum of the product state elements is trivially equivariant with respect to similar graphs. If we have $n=2$ final qubit states, the product state is
@@
for commutative spaces where $v_{i}^{j}\in\mathbb{C}$ and $\{1,2\},\{2,1\}$ should be regarded as ordered sets, which again shows the sum of the state elements remaining unchanged when the qubit states switch positions in the product. We now assume the statement is true for $n=N$ final qubit states and proceed to show the $N+1$ case is true. The quantum product state over $N$ elements becomes
@@
which we assume to be permutation-equivariant over the sum of its elements. We can rewrite the form of this state as
@@
where $A_{j}$ defines the $2^{N}$ terms in the final product state. Replacing the $i+1$th entry of the Kronecker product above with a new $N+1$th state, we have
@@
When this occurs, this new state consisting of $2^{N+1}$ elements with the $N+1$ state in the $i+1$th entry of the product can be written in terms of the old state with groupings of the new elements in $2^{N+1-i}$ batches of $2^{i}$ elements, i.e.,
@@
which, when summed, becomes
@@
However, the $i+1$th entry is arbitrary, and, due to the summation permutation equivariance of the initial state $\bigotimes_{i=1}^{N}|\psi_{i}\rangle$, the sum $\sum_{j=1}^{2^{N}}A_{j}$ is equivariant, in fact invariant, under all reorderings of the elements $|\psi_{i}\rangle$ in the product $\bigotimes_{i=1}^{N}|\psi_{i}\rangle$. Therefore, we conclude $\bigotimes_{i=1}^{N+1}|\psi_{i}\rangle$ is permutation-equivariant with respect to the sum of its elements. ∎
To show a simple illustration of why (13) is true, let us take two initial states and see what happens when we insert a new state between them, i.e., in the $2$nd entry in the product. This should lead to $42^{2+1-1}=2^{2}=$ 4 groupings of $22^{1}=$ 2 elements. To begin, we have
@@
and when we insert the new third state in the $1$st entry of the product above, we have